# The generalized Hierarchical Gaussian Filter

### v2

@@ Abstract
Hierarchical Bayesian models of perception and learning feature prominently in contemporary cognitive neuroscience where, for example, they inform computational concepts of mental disorders. This includes predictive coding and hierarchical Gaussian filtering (HGF), which differ in the nature of hierarchical representations. Predictive coding assumes that higher levels in a given hierarchy influence the state (value) of lower levels. In HGF, however, higher levels determine the rate of change at lower levels. Here, we extend the space of generative models underlying HGF to include a form of nonlinear hierarchical coupling between state values akin to predictive coding and artificial neural networks in general. We derive the update equations corresponding to this generalization of HGF and conceptualize them as connecting a network of (belief) nodes where parent nodes either predict the state of child nodes or their rate of change. This enables us to (1) create modular architectures with generic computational steps in each node of the network, and (2) disclose the hierarchical message passing implied by generalized HGF models and to compare this to comparable schemes under predictive coding. We find that the algorithmic architecture instantiated by the generalized HGF is largely compatible with that of predictive coding but extends it with some unique predictions which arise from precision and volatility related computations. Our developments enable highly flexible implementations of hierarchical Bayesian models for empirical data analysis and are available as open source software.
Keywords: hierarchical Gaussian filter, HGF, predictive coding, perceptual inference, neuromodelling, computational psychiatry
@@ 1 Introduction
1.1 Overview
In order successfully to navigate complex environments, biological agents need to combine noisy sensory inputs with prior knowledge to infer on the true, but hidden, state of the world. Hierarchical Bayesian models such as predictive coding (Rao & Ballard, 1999; Friston, 2005) have become popular tools to understand brain functions which enable humans to form beliefs based on sparse and ambiguous sensory inputs. These models assume that the brain constructs a hierarchy of beliefs, with higher level beliefs relating to increasingly abstract features of the world. For example, hierarchical Gaussian filtering (HGF, Mathys et al., 2011, 2014a) models a hierarchy in which higher levels encode the rate of change in lower level features, enabling agents adaptively to scale their learning behaviour in response to changes in the stability of their environment. Here, we extend this framework to a different type of hierarchy, where higher-level beliefs influence lower-level beliefs via their expectation. In other words, higher-level beliefs can here also determine the value of lower-level beliefs instead of only their rate of change. This extension significantly widens hierarchical Gaussian filtering’s application scope and enables a direct comparison of the implied message-passing scheme with, for example, predictive coding models. Moreover, the new derivations equip HGF with a modular architecture which allows for more versatile and user-friendly implementations of HGF models.
@@ 1.2 Hierarchical Bayesian models of perception and learning
Bayesian perspectives on perception have proposed that our brain inverts a hierarchical generative model to infer the causes of its sensory inputs and predict future events (Dayan et al., 1995; Rao & Ballard, 1999; Friston, 2010; Doya et al., 2011; Helmholtz, 1860). Under this “Bayesian brain” view, perception corresponds to integrating expectations (prior beliefs about hidden states of the world) with incoming sensory information to yield a posterior belief, while learning refers to the updating of beliefs about the model’s parameters, which takes place more slowly, as experience accumulates. Formally, beliefs are modelled as probability distributions, such that the width of the distribution reflects the uncertainty (inverse precision) associated with that belief. Humans have been shown to take uncertainty into account when combining different sources of information in a manner that conforms to the statistical optimum as prescribed by Bayes’ rule (Ernst & Banks, 2002; Angelaki et al., 2009).
@@
Because sensory signals are generated by interacting causes in the external environment that span multiple spatial and temporal scales, the brain in its Helmholtzian description is assumed to reflect this hierarchy of causes in a correspondingly hierarchical generative model. The Bayesian inversion of this generative model results in a hierarchy of beliefs, where higher levels encode beliefs about increasingly abstract, general, and stable features of the environment. These higher-level beliefs serve as priors for the inference on lower levels. Specifically, at each level of the hierarchy, belief updates serve to reconcile predictions (priors) from higher levels with the actual input (likelihood) from lower levels. Furthermore, under fairly general assumptions (i.e., for all probability distributions from the exponential family, Mathys, 2016; Mathys & Weber, 2020), these belief updates rest on (precision-weighted) prediction errors (PEs), i.e., the (weighted) mismatch between the model’s predictions and the actual input (Friston, 2010).
@@
Popular hierarchical Bayesian models of perception and learning that are built on these ideas are predictive coding (Rao & Ballard, 1999; Friston, 2005) and hierarchical Gaussian filtering (Mathys et al., 2011, 2014a). In these models, estimates of uncertainty are central: the impact of prediction errors on belief updates depends on a precision ratio, which relates the precision of the prior to that of the observation, thus scaling the relative impact that new information has on belief updates. Put simply, mismatches (PEs) elicit stronger belief updates if the prediction about the input (likelihood) is precise, relative to the belief in the current estimate (prior). This form of adaptive scaling, a key element of healthy inference and learning, has also been proposed to lie at the heart of perceptual disturbances observed in mental disorders. For example, an imbalance between the influence of expectations and sensory inputs has featured prominently in attempts to explain the emergence of positive symptoms in schizophrenia, such as hallucinations and delusions (Stephan et al., 2006; Fletcher & Frith, 2009; Corlett et al., 2009, 2011; Adams et al., 2013; Friston et al., 2016; Sterzer et al., 2018).
@@
The HGF111we use “the HGF” as a shorthand for “hierarchical Gaussian filtering models of various configurations” has been particularly useful in this context, as it can be fit to participants’ empirically observed behaviour or physiology, and thereby used to infer individual trajectories of precision-weighted PEs and predictions from data. By formulating a response model that links trial-wise perceptual quantities (such as predictions and PEs) to measured quantities (such as choices, reaction times, eye movements, evoked response amplitude in EEG, etc.), the HGF can quantify individual differences in inference and learning in terms of model parameters that encode prior beliefs about higher-order structure in the environment (de Berker et al., 2016; Lawson et al., 2017; Powers et al., 2017; Siegel et al., 2020; Sevgi et al., 2020; Henco et al., 2020; Rossi-Goldthorpe et al., 2021; Suthaharan et al., 2021; Kafadar et al., 2022; Sapey-Triomphe et al., 2022; Fromm et al., 2023; Drusko et al., 2023). Such a mechanistic characterization of inter-subject variability is of particular interest for fields like Computational Psychiatry, because such differences may explain the heterogeneous nature of psychiatric diseases, and form a basis for mapping them out in a continuous conceptual space or dividing them into more homogeneous subgroups (Stephan & Mathys, 2014; Mathys, 2016).
@@ 1.3 Volatility-coupling, value-coupling, and noise-coupling
The type of belief hierarchy modelled by any particular approach depends on the nature of the generative model. The HGF assumes a particular form of generative model, where hidden states of the world evolve as coupled random walks in time. In current HGF models, the mean (value) of the higher level determines the variance (step size) of the lower level’s random walk. In other words, higher levels encode the volatility (or inverse stability) of lower levels (we will call this volatility coupling). This is motivated by the observation that learners must take into account different sources of uncertainty in their belief updates, one of which is the current level of stability in the world: if the world is currently changing (volatile), the agent needs to learn faster. Accordingly, in the HGF, subjective [^2] Note that the word subjective does not imply conscious accessibility here - it only differentiates the inferred quantities from the “true“ values as produced by the environment or the generative model. estimates of increased environmental volatility directly influence the uncertainty associated with lower level beliefs, leading to faster belief updates on the lower level. Previous work has shown that human learners indeed adjust their learning rate according to experimentally manipulated levels of volatility (Behrens et al., 2007).
@@
By contrast, predictive coding models typically focus on hierarchies in which higher levels predict the value of lower levels, or, in other words, the mean of the probability distribution that represents the lower-level belief, or in other words, the expectation of the value. We will refer to these hierarchies as implementing value coupling. This type of hierarchy is useful for understanding how beliefs about lower-level features depend on higher-level beliefs – for example, how the perceived brightness of a patch in an image depends on the context (objects, shadows) in which that patch is presented (Rao & Ballard, 1999; Adelson, 2005). It is worth noting that the type of hierarchical coupling in predictive coding is often used in theoretical treatments of hierarchical Bayesian modelling while the HGF offers a flexibly applicable implementation that is being widely used for empirical data analysis.
@@
In a noteworthy exception to this, Kanai et al. (2015) presented a predictive coding model where higher levels encode the (spatial) precision of lower levels in a static environment. This is different from volatility coupling, where higher levels are concerned with the rate of change on lower levels, but captures a second source of uncertainty in beliefs about hidden states: the level of noise or reliability of the sensory input.
Relatedly, in the learning and decision-making literature, two classes of models separately deal with the estimation of process noise (volatility, Behrens et al., 2007; Mathys et al., 2011; Piray & Daw, 2020) and observation noise (stochasticity, Lee et al., 2020; Nassar et al., 2010), with recent attempts to capture both sources of uncertainty in a joint model (Piray & Daw, 2021).
In this work, we show that the HGF can be extended to encompass both volatility and value coupling both at the level of hidden states, and at the observation stage. Importantly, coupling types that link levels via lower-level belief precision include volatility coupling (between hidden states) as well as a principled way of modelling inference on observation noise: via noise coupling at the level of observations. Together with the different types of value coupling we will introduce, the generalised HGF thus provides a very general modelling framework with a wide range of potential applications.
1.4 Contribution of current work
In this technical note, we extend the generative model of the HGF to consider hierarchical value coupling alongside volatility coupling. Based on the work in Weber (2020), we (1) derive simple, efficient one-step belief update equations for linear and non-linear value coupling, and (2) conceptualize these equations, together with their volatility coupling counterparts, as a network of interacting nodes which can be implemented in a modular architecture.
In brief, we show that:
  1. 1.
    The HGF provides a very general modelling framework that encompasses multiple types of interactions between states in the world - where higher-level hidden states determine the value, the rate of change (volatility), or even the level of noise in lower-level states or observations.
  2. 2.
    The explicit treatment of volatility estimation in the HGF allows for an implementation that comprises both global control of volatility-related precision (implemented by a global high-level volatility belief that affects multiple low-level states), and local or distributed volatility estimation, enabling modality-specific modulation of learning rates.
  3. 3.
    The message passing scheme for value coupling in the HGF is almost equivalent to recently proposed predictive coding architectures, apart from small, but interesting differences. These differences relate on one hand to the discrete nature of the updates in the HGF, and, on the other hand, to the volatility-related updates of belief uncertainty.
2 Introducing value coupling
@@
In the original formulation (Mathys et al., 2011, 2014a), coupling between environmental states at different hierarchical levels was implemented in the form of volatility coupling, where the step size $\vartheta$ (or rate of change/volatility) of a state $x_{a}$ varies a function of a higher-level state $x_{\check{a}}$:
@@
$x_{a}^{(k)}\sim\mathcal{N}\left(x_{a}^{(k-1)}+\rho_{a}+f\left(x_{b}^{(k)}\right),\,\vartheta\right),$ (3)
@@
$f\left(x_{b}^{(k)}\right)=\alpha_{b,a}\,g_{b,a}\left(x_{b}^{(k)}\right)$ (4)
is the coupling function with parameter $\alpha_{b,a}$ (scaling the impact of value parent $x_{b}$ on $x_{a}$). The state’s mean now evolves as a function of its previous value $x_{a}^{(k-1)}$, a drift parameter $\rho_{a}$, and a function of state $x_{b}^{(k)}$. Crucially, as we will show below, the HGF can deal with both linear and nonlinear transformation functions $g$, as long as the function $g$ is twice continuously differentiable almost everywhere. States $x_{b}$ and $x_{a}$ interact via value coupling.
@@
$x_{a}^{(k)}\sim\mathcal{N}\left(\lambda_{a}x_{a}^{(k-1)}+\rho_{a}+f\left(x_{b}^{(k)}\right),\,\vartheta\right).$ (5)
@@
$M_{a}=\frac{\rho_{a}+f\left(x_{b}^{(k)}\right)}{1-\lambda_{a}},$ (6)
and the value of $\lambda$ will determine the speed with which it does so. In other words, the meaning of $\rho$ and $x_{b}$ for the evolution of $x_{a}$ depends on $\lambda$. The difference between the cases $\lambda=1$ and $\lambda<1$ is illustrated in Figure 1, and the derivation based on a first-order autoregressive process is given in Appendix LABEL:sec:appautoreg). Finally, if $\lambda=0$, the state does not perform a Gaussian random walk around its own previous mean anymore, instead, its values are normally distributed around a constant or the value of a parent state (reminiscent of an observation of the parent with Gaussian noise).
Importantly, different forms of coupling can be present at the same time: A state $x_{a}$ can have both a volatility parent $x_{\check{a}}$ (generating changes in its rate of change) and a value parent $x_{b}$ (generating changes in its mean value). It can also have a drift parameter $\rho_{a}$ which is a constant influencing its mean – equivalent to its tonic volatility parameter $\omega_{a}$ which determines its step size in the absence of a phasic volatility influence. Finally, we allow for inputs to arrive at irregular intervals; therefore, we multiply the total variance of the random walk and the total mean drift by the time $t^{(k)}$ that has passed since the arrival of the previous sensory input at index $k-1$ (Mathys et al., 2014a). Together with suitably chosen priors on parameters and initial states (see Mathys et al., 2014a), the following equation forms the generative model for a state $x_{a}$ with both volatility and value parent:
@@
Because a given state can also be parent to multiple child states at the same time, these extensions allow us to model fairly complex networks of interacting states of the world. In Figure 2 we have drawn an example setup with 11 different environmental states and two outcomes. For this example, and together with priors on parameters and initial states (see Mathys et al., 2014a), the following equations describe the generative model (for simplicity, the example uses linear value coupling, no tonic drifts ($\rho=0$, $\lambda=1$), and inputs at regular intervals, i.e., $t^{(k)}\equiv 1\;\forall\,k$ ):
@@
Using this example network, we introduce two general motifs. First, all states that are value parents of other states (or outcomes) by default have their own volatility parent (and volatility parents therefore share the index with their child node, for example, states $x_{a}$ and $x_{\check{a}}$). Even if in practice, many environmental states might have constant volatility, from the perspective of the agent, it makes sense to a-priori allow for phasic changes in volatility. From a modelling perspective, these volatility parents could be removed in scenarios with constant volatility. On the other hand, more than one volatility levels may sometimes be required.
Second, states that are volatility parents to other states can either have a value parent (as states $x_{\check{a}}$ and $x_{\check{d}}$), or no parents (as states $x_{\check{c}}$$x_{\check{e}}$ and $x_{\check{f}}$). This is because in practice, volatility parents of volatility parents are rarely required. Instead, we suggest that value parents of volatility states are more useful. In particular, these can be used to model “global” or shared volatility states that affect multiple lower-level volatility beliefs (such as state $x_{f}$ in this example, which influences volatility beliefs about states $x_{a}$ and $x_{d}$), separately from “local” volatility states that only affect speed of change in a single lower-level state (such as the volatility states $x_{\check{c}}$$x_{\check{e}}$ and $x_{\check{f}}$).
@@
In the approximate inversion of the generative model for volatility coupling, Mathys et al. (2011) derived a set of single-step update equations that represent the approximately Bayes-optimal changes in these beliefs in response to incoming stimuli. Repeating this derivation for the case of value coupling similarly leads us to simple one-step equations for updating beliefs about states that serve as value parents (for the full derivation of these equations, cf. Appendix 6.1). Assuming state $x_{b}$ is a value parent to state $x_{a}$ with a coupling strength $\alpha_{b,a}$, then the new posterior belief about state $x_{b}$ after observing a new input at time step $k$ is given by:
@@
To get an intuition for these update equations in the case of nonlinear value coupling, let us consider an example where state $x_{a}$ acts as a rectified linear unit (ReLU), that is, we choose the function $g$ as the rectifier function, a popular activation function for deep neural networks (Hahnloser et al., 2000; Lecun et al., 2015):

### v3

@@ Abstract
Hierarchical Bayesian models of perception and learning feature prominently in contemporary cognitive neuroscience where, for example, they inform computational concepts of mental disorders. This includes predictive coding and hierarchical Gaussian filtering (HGF), which differ in the nature of hierarchical representations. In this work, we present a new class of artificial neural networks that unifies computational principles of PC and HGFs. We extend the space of generative models underlying HGF to include a form of nonlinear hierarchical coupling between state values akin to predictive coding and artificial neural networks in general. We derive the update equations corresponding to this generalization of HGF and conceptualize them as connecting a network of (belief) nodes where parent nodes either predict the state of child nodes or their rate of change. This enables us to (1) create modular architectures with generic computational steps in each node of the network, and (2) disclose the hierarchical message passing implied by generalized HGF models and to compare this to comparable schemes under predictive coding. The practical advances of this work are twofold: on the one hand, our extension allows for a modular construction of ANNs of arbitrarily complex hierarchical structure under the general principles of HGF. On the other hand, by providing a highly flexible implementation of hierarchical Bayesian models available as open source software, it enables new types of empirical data analysis in computational psychiatry.
Keywords: hierarchical Gaussian filter, HGF, predictive coding, artificial neural networks, perceptual inference, neuromodelling, computational psychiatry
@@ 1.1 Hierarchical Bayesian models of perception and learning
Bayesian perspectives on perception have proposed that our brain inverts a hierarchical generative model to infer the causes of its sensory inputs and predict future events (Dayan1995; Rao1999; Friston2010; Doya2011). Under this “Bayesian brain” view, perception corresponds to integrating expectations (prior beliefs about hidden states of the world) with incoming sensory information to yield a posterior belief, while learning refers to the updating of beliefs about the model’s parameters, which takes place more slowly, as experience accumulates. Formally, beliefs are modelled as probability distributions, such that the width of the distribution reflects the uncertainty (inverse precision) associated with that belief. Humans have been shown to take uncertainty into account when combining different sources of information in a manner that conforms to the statistical optimum as prescribed by Bayes’ rule (ernst2002; Angelaki2009).
@@
Because sensory signals are generated by interacting causes in the external environment that span multiple spatial and temporal scales, the brain is assumed to reflect this hierarchy of causes in a correspondingly hierarchical generative model. The Bayesian inversion of this generative model results in a hierarchy of beliefs, where higher levels encode beliefs about increasingly abstract, general, and stable features of the environment. These higher-level beliefs serve as priors for the inference on lower levels. Specifically, at each level of the hierarchy, belief updates serve to reconcile predictions (priors) from higher levels with the actual input (likelihood) from lower levels. Furthermore, under fairly general assumptions (i.e., for all probability distributions from the exponential family, mathys2016; mathys_weber_2020), these belief updates rest on (precision-weighted) prediction errors (PEs), i.e., the (weighted) mismatch between the model’s predictions and the actual input (Friston2010).
@@
Popular hierarchical Bayesian models of perception and learning that are built on these ideas are predictive coding (Rao1999; Friston2005) and hierarchical Gaussian filtering (mathys2011; mathys2014). In these models, estimates of uncertainty are central: the impact of prediction errors on belief updates depends on a precision ratio, which relates the precision of the prior to that of the observation, thus scaling the relative impact that new information has on belief updates. Put simply, mismatches (PEs) elicit stronger belief updates if the prediction about the input (likelihood) is precise, relative to the belief in the current estimate (prior). This form of adaptive scaling, a key element of healthy inference and learning, has also been proposed to lie at the heart of perceptual disturbances observed in mental disorders. For example, an imbalance between the influence of expectations and sensory inputs has featured prominently in attempts to explain the emergence of positive symptoms in schizophrenia, such as hallucinations and delusions (Stephan2006; Fletcher2009; Corlett2009; Corlett2011; Adams2013psycho; Friston2016; sterzer2018; corlett2019hallu).
@@
The HGF111we use “the HGF” as a shorthand for “hierarchical Gaussian filtering models of various configurations” has been particularly useful in this context, as it can be fit to participants’ empirically observed behaviour or physiology, and thereby used to infer individual trajectories of precision-weighted PEs and predictions from data. By formulating a response model that links trial-wise perceptual quantities (such as predictions and PEs) to measured quantities (such as choices, reaction times, eye movements, evoked response amplitude in EEG, etc.), the HGF can quantify individual differences in inference and learning in terms of model parameters that encode prior beliefs about higher-order structure in the environment (de_berker_computations_2016; lawson_adults_2017; powers_pavlovian_2017; siegel_computational_2020; sevgi_social_2020; henco_bayesian_2020; rossi-goldthorpe_paranoia_2021; suthaharan_paranoia_2021; kafadar_conditioned_2022; sapey-triomphe_associative_2022; fromm_belief_2023; drusko_novel_2023). Such a mechanistic characterization of inter-subject variability is of particular interest for fields like Computational Psychiatry, because such differences may explain the heterogeneous nature of psychiatric diseases, and form a basis for mapping them out in a continuous conceptual space or dividing them into more homogeneous subgroups (stephan2014; mathys2016).
@@ 1.2 Volatility-coupling, value-coupling, and noise-coupling
The type of belief hierarchy modelled by any particular approach depends on the nature of the generative model. The HGF assumes a particular form of generative model, where hidden states of the world evolve as coupled random walks in time. In current HGF models, the mean (value) of the higher level determines the variance (step size) of the lower level’s random walk. In other words, higher levels encode the volatility (or inverse stability) of lower levels (we will call this volatility coupling). This is motivated by the observation that learners must take into account different sources of uncertainty in their belief updates, one of which is the current level of stability in the world: if the world is currently changing (volatile), the agent needs to learn faster. Accordingly, in the HGF, subjective [^2] Note that the word subjective does not imply conscious accessibility here - it only differentiates the inferred quantities from the “true“ values as produced by the environment or the generative model. estimates of increased environmental volatility directly influence the uncertainty associated with lower level beliefs, leading to faster belief updates on the lower level. Previous work has shown that human learners indeed adjust their learning rate according to experimentally manipulated levels of volatility (behrens2007).
@@
By contrast, predictive coding models typically focus on hierarchies in which higher levels predict the value of lower levels. This is implemented as the mean of the probability distribution that represents the lower-level belief, or in other words, the expectation of the value. We will refer to these hierarchies as implementing value coupling. This type of hierarchy is useful for understanding how beliefs about lower-level features depend on higher-level beliefs – for example, how the perceived brightness of a patch in an image depends on the context (objects, shadows) in which that patch is presented (Rao1999; adelson2005). Modern predictive coding neural networks are powerful machine learning tools (millidge2022; song2020; whittington2017). While value coupling - the type of hierarchical coupling in predictive coding - is also often used in theoretical treatments of hierarchical Bayesian modelling, the HGF, in contrast, offers a flexibly applicable implementation of hierarchical beliefs connected by volatility coupling that is being widely used for empirical data analysis.
@@
In a noteworthy exception to the typical value hierarchies in predictive coding, kanai2015 presented a model where higher levels encode the (spatial) precision of lower levels in a static environment. This is different from volatility coupling, where higher levels are concerned with the rate of change on lower levels, but captures a second source of uncertainty in beliefs about hidden states: the level of noise or reliability of the sensory input. Relatedly, in the learning and decision-making literature, two classes of models separately deal with the estimation of process noise (volatility, behrens2007; mathys2011; piray2020a) and observation noise (stochasticity, lee2020; nassar2010), with recent attempts to capture both sources of uncertainty in a joint model (piray2020b). Importantly, both volatility and stochasticity estimation can be captured as higher-level belief nodes coupled to the precision of lower-level beliefs.
Here, we show that the HGF can be extended to encompass all of these architectures: precision coupling between hidden states (volatility coupling) and at the observation stage (noise coupling) and, importantly, linear and nonlinear forms of value coupling. Our extension thus links HGF with predictive coding and also includes a principled way of modelling inference on observation noise.
The generalised HGF represents a new class of artificial neural networks that unifies computational principles of predictive coding and HGF. Practically, it allows for modular construction of ANNs with complex hierarchical structure, and it provides a flexible tool for empirical data analysis in computational psychiatry via implementation in open software. These theoretical and practical advances will be unpacked in the following sections.
2 Theory: Introducing value coupling
@@
In the original formulation (mathys2011; mathys2014), coupling between environmental states at different hierarchical levels was implemented in the form of volatility coupling, where the step size $\vartheta$ (or rate of change/volatility) of a state $x_{a}$ varies a function of a higher-level state $x_{\check{a}}$:
@@
$x_{a}^{(k)}\sim\mathcal{N}\left(x_{a}^{(k-1)}+f\left(x_{b}^{(k)}\right),\,\vartheta\right),$ (3)
@@
$f\left(x_{b}^{(k)}\right)=\alpha_{b,a}\,g_{b,a}\left(x_{b}^{(k)}\right)+\rho_{a}$ (4)
is the coupling function with parameter $\alpha_{b,a}$ (scaling the impact of value parent $x_{b}$ on $x_{a}$), the constant drift parameter $\rho_{a}$, and the function $g$. The state’s mean now evolves as a function of its previous value $x_{a}^{(k-1)}$, and a function of state $x_{b}^{(k)}$. Crucially, as we will show below, the HGF can deal with both linear and nonlinear transformation functions $g$, as long as the function $g$ is twice continuously differentiable almost everywhere. States $x_{b}$ and $x_{a}$ interact via value coupling.
@@
$x_{a}^{(k)}\sim\mathcal{N}\left(\lambda_{a}x_{a}^{(k-1)}+f\left(x_{b}^{(k)}\right),\,\vartheta\right).$ (5)
@@
$M_{a}=\frac{f\left(x_{b}^{(k)}\right)}{1-\lambda_{a}},$ (6)
and the value of $\lambda$ will determine the speed with which it does so. In other words, the meaning of $\rho$ and $x_{b}$ for the evolution of $x_{a}$ depends on $\lambda$. The difference between the cases $\lambda=1$ and $\lambda<1$ is illustrated in Figure 1. Finally, if $\lambda=0$, the state does not perform a Gaussian random walk around its own previous mean anymore, instead, its values are normally distributed around a constant or the value of a parent state (reminiscent of an observation of the parent with Gaussian noise).
Importantly, different forms of coupling can be present at the same time: A state $x_{a}$ can have both a volatility parent $x_{\check{a}}$ (generating changes in its rate of change) and a value parent $x_{b}$ (generating changes in its mean value). It can also have a drift parameter $\rho_{a}$ which is a constant influencing its mean – equivalent to its tonic volatility parameter $\omega_{a}$ which determines its step size in the absence of a phasic volatility influence. Finally, we allow for inputs to arrive at irregular intervals; therefore, we multiply the total variance of the random walk and the total mean drift by the time $t^{(k)}$ that has passed since the arrival of the previous sensory input at index k1k-1 (mathys2014). Together with suitably chosen priors on parameters and initial states (see mathys2014), the following equation forms the generative model for a state $x_{a}$ with both volatility and value parent:
@@
Because a given state can also be parent to multiple child states at the same time, these extensions allow us to model fairly complex networks of interacting states of the world. In Figure 2 we have drawn an example setup with 11 different environmental states and two outcomes. For this example, and together with priors on parameters and initial states (see mathys2014), the following equations describe the generative model (for simplicity, the example uses linear value coupling, no tonic drifts ($\rho=0$, $\lambda=1$), and inputs at regular intervals, i.e., $t^{(k)}\equiv 1\;\forall\,k$ ):
@@
Using this example network, we introduce two general motifs. First, all states that are value parents of other states (or outcomes) by default have their own volatility parent (and volatility parents therefore share the index with their child node, for example, states $x_{a}$ and $x_{\check{a}}$). Even if in practice, many environmental states might have constant volatility, from the perspective of the agent, it makes sense to a-priori allow for phasic changes in volatility. From a modelling perspective, these volatility parents could be removed in scenarios with constant volatility.
Second, states that are volatility parents to other states can either have a value parent (as states $x_{\check{a}}$ and $x_{\check{d}}$), or no parents (as states $x_{\check{c}}$, $x_{\check{e}}$ and $x_{\check{f}}$). This is because in practice, volatility parents of volatility parents are rarely required. Instead, we suggest that value parents of volatility states are more useful. In particular, these can be used to model “global” or shared volatility states that affect multiple lower-level volatility beliefs (such as state $x_{f}$ in this example, which influences volatility beliefs about states $x_{a}$ and $x_{d}$), separately from “local” volatility states that only affect speed of change in a single lower-level state (such as the volatility states $x_{\check{c}}$, $x_{\check{e}}$ and $x_{\check{f}}$).
@@
In the approximate inversion of the generative model for volatility coupling, mathys2011 derived a set of single-step update equations that represent the approximately Bayes-optimal changes in these beliefs in response to incoming stimuli. Repeating this derivation for the case of value coupling similarly leads us to simple one-step equations for updating beliefs about states that serve as value parents (for the full derivation of these equations, cf. Appendix 6.1). Assuming state $x_{b}$ is a value parent to state $x_{a}$ with a coupling strength $\alpha_{b,a}$, then the new posterior belief about state $x_{b}$ after observing a new input at time step $k$ is given by:
@@
To get an intuition for these update equations in the case of nonlinear value coupling, let us consider an example where state $x_{a}$ acts as a rectified linear unit (ReLU), that is, we choose the function $g$ as the rectifier function, a popular activation function for deep neural networks (Hahnloser2000; Lecun2015):